OpenStudy (anonymous):

can someone help me on this !!! -3y=x-6 i have to find x and y so i can put it in a graph but i don't understand it so someone PLEASE help me(:

OpenStudy (anonymous):

x=6−3y :)

OpenStudy (texaschic101):

are you looking for the x and y intercepts ?

OpenStudy (anonymous):

ya but i took to long to answer and i got it wrong.. :(

OpenStudy (texaschic101):

do you want to know how to do it ? For future problems

OpenStudy (anonymous):

So do you need the graph now or is it to late

OpenStudy (anonymous):

i have another one you could help me on if you want..

OpenStudy (anonymous):

Sorry i couldnt catch it on time!

OpenStudy (texaschic101):

ask away

OpenStudy (anonymous):

Whats the problem and do you need the graph or just the answer?

OpenStudy (anonymous):

ill let @texaschic101 help you good luck

OpenStudy (anonymous):

@Monica75 both but okay thanks anyways (:

OpenStudy (anonymous):

-5x-3y=-7

OpenStudy (anonymous):

Answer x=7\5−3y\5

OpenStudy (anonymous):

:) DID THAT HELP

OpenStudy (anonymous):

yes it did thanks (:

OpenStudy (anonymous):

WELCOME :) :) anytime

OpenStudy (anonymous):

are you from new mexico?

OpenStudy (texaschic101):

-5x - 3y = -7 Lets put this in y = mx + b form -3y = 5x - 7 y = -5/3x + 7/3 The y intercept in y = mx + b form is in the b position...so the y intercept is (7/3)or (0,7/3)..for graphing purposes, 7/3 = 2 1/3 To find the x intercept, sub in 0 for y in your equation -5x - 3y = -7 -5x - 3(0) = -7 -5x = -7 x = 7/5....the x intercept is 7/5 or (7/5,0)...for graphing purposes, 7/5 = 1 2/5 In y = mx + b form, the slope is in the m position, therefore, the slope is -5/3 so we have : y intercept = (0,7/3) x intercept = (7/5,0) slope = -5/3

OpenStudy (anonymous):

no but i am half mexican lol